OpenStudy (anonymous):

find the following sum: 7z/15x+4z/9x

OpenStudy (matt101):

Find the common denominator, which is 45x. Then you multiply the numerator and denominator of both terms by the appropriate number to get 45x as the denominator. (3*7z)/(3*15x) + (5*4z)/(5*9x) = 21z/45x + 20z/45x = 41z/45x

OpenStudy (anonymous):

\[\frac{41 x z}{45} \] http://www.wolframalpha.com/input/?i=simplify7+z%2F15+x+%2B+4+z%2F9+x&t=sftb01

OpenStudy (matt101):

That's true, but I think they meant to have x in the denominator. Mind you I guess my solution is just as ambiguous...ah the joy of brackets...

OpenStudy (anonymous):

Starting today, I will warn them that the normal precedence rules will be applied if no parentheses or brackets are applied.